# A Survey on Human-AI Teaming with Large Pre-Trained Models

### v1

@@ 5 Applications
This section highlights the role of Human-AI (HAI) Teaming and Large Pre-trained Models (LPtM) in revolutionizing fields such as healthcare, autonomous vehicle domain, surveillance, gaming, education, and accessibility. It underscores the synergy between human expertise and AI’s analytical power, emphasizing the importance of ethical AI development and user-centric design in maximizing benefits and inclusivity within society.
5.1 Healthcare
@@
In the healthcare sector, the symbiosis of human-AI collaboration, especially in enhancing patient diagnosis accuracy and fostering trust, underscores AI’s transformative influence [63, 11, 104]. Advances such as the application of AI in breast cancer detection have markedly boosted diagnostic efficacy, sometimes even outperforming human experts, showcasing the remarkable potential of AI in healthcare [102]. These developments necessitate reevaluating AI’s role in patient care, addressing ethical considerations and practical challenges [16]. Research focusing on healthcare practitioners’ views within the US highlights the critical factors of training, trustworthiness, and perceived AI risks, which are pivotal in its acceptance and application in clinical environments [27]. Additionally, improvements in AI-powered medical image retrieval systems have been shown to augment both diagnostic capabilities and user trust, advocating for integrated systems that blend human insight with algorithmic accuracy [17]. This partnership between AI technologies and healthcare professionals is vital for enhancing patient care outcomes and alleviating professional strain, albeit confronted by hurdles related to training and the accountability of AI systems.
@@
Integrating Large Pre-trained Models such as ChatGPT into healthcare has shown promise in enhancing medical education, patient communication, and clinical decision-making. Lyu et al. [99] demonstrated ChatGPT’s ability to simplify radiology reports for patients and healthcare providers, highlighting the potential for improved clinical education despite noting areas for accuracy and specificity enhancements. Kung et al. [86] found that ChatGPT scored at or near the passing threshold on the USMLE, suggesting its capability to support medical education and hinting at future roles in clinical settings. Johnson et al. [71] further validated ChatGPT’s usefulness by showing its generally accurate responses to a range of medical questions, though recognizing the need for further refinement.
5.2 Autonomous Vehicles
@@
Advancements in autonomous driving promise safer, AI-enhanced transportation, with significant milestones achieved in vehicle sensing and decision-making [4]. Key to these advancements is understanding Human-AI collaboration, particularly during critical transitions between automated and manual driving modes, as explored by Lv et al. [97]. This research highlights the importance of seamless transitions for safety, showing improvements in control and stability with a new hybrid driving model. Additionally, establishing trust in AI through transparent decision-making processes under uncertainty is crucial, as Jianlong et al. [180] discuss with a novel framework aimed at enhancing Human-AI collaboration. Fuchs et al. [48] further examine the managerial aspects of task delegation between humans and AI, proposing reinforcement learning to optimize collaboration and improve system performance.
@@
Building on foundational studies, recent explorations into large language models (LLMs) signal a transformative era in autonomous driving technologies. Yang et al.[174] underscore the importance of LLMs for interpreting user commands in complex scenarios, emphasizing the necessity of model quality and prompt design. Park et al. [124] introduce VLAAD, a multi-modal LLM that demonstrates proficient interpretive capabilities across varied driving situations, showcasing the potential of natural language models in enhancing decision-making processes. Cui et al. [35] propose a framework that integrates LLMs for improved decision-making in autonomous vehicles, aiming for personalized assistance and efficient operation. Wen et al. [167] highlight the role of Visual Language Models (VLMs) like GPT-4V in scene understanding and decision-making, noting superior performance but also pointing out existing challenges. Additionally, Li et al. [91] tackle the challenge of generating realistic multi-view video data for autonomous driving, further supporting the role of LLMs in advancing the field. Together, these studies underscore the transformative potential of LLMs in advancing autonomous vehicle technologies, from improving interpretive capabilities to enhancing decision-making processes and data generation.
5.3 Surveillance and Security
@@
The advancements in vision-based AI and IoT-connected cameras have heightened the potential for AI in cybersecurity [125], including in Cyber Security Incident Response Teams (CSIRTs) [78]. Determining the optimal level of AI autonomy in these settings is critical. A study by Hauptman et al. [60] explored this through a mixed-method approach, assessing AI’s role across the CSIRT cycle—preparation, identification, containment, eradication, and recovery. Surveying 103 cybersecurity professionals and conducting interviews with 22, the study found that higher AI autonomy is preferred during the identification phase, while lower autonomy suits high-risk phases better. This suggests that adaptive AI, with proper team training, can integrate effectively into CSIRTs, enhancing teamwork and response efficiency. Furthermore, the real-time update preference and successful integration of human and machine intelligence in crowd-powered camera sensing, as noted by Guo et al. [55], underscore the potential for enhancing cybersecurity measures through improved situational awareness and resource management.
@@
Moreover, the utilization of LLMs for video surveillance analysis, as proposed by Chen et al. [23], with the VideoLLM framework, demonstrates the potential for transferring LLMs’ understanding and reasoning capabilities to video understanding tasks, offering a unified approach for various video-based security challenges. Additionally, Jain’s [68] insights into integrating ChatGPT with video security underscore the transformative impact of foundation models on physical security systems, enhancing operational efficiency and surfacing new insights for businesses.
5.4 Games
@@
5.4.1 LLM-Based AI as Players
Artificial intelligence has long been utilized in video games, from serving as basic adversaries in titles like Mario Kart [145] to mastering complex games such as chess, demonstrating AI’s algorithmic prowess. Yet, these AIs often fell short as complete substitutes for human players due to their lack of flexibility and limited language processing capabilities. The advent of large language models (LLMs) has shifted this landscape, offering AIs that can engage with humans in more expressive ways, suitable across a broader spectrum of games.
Research into LLMs as game players, while in its nascent stages, has yielded intriguing insights. For instance, Frans et al. [46] showcase "AI Charades" using GPT-2 to illustrate LLMs’ potential in games requiring deep language interaction, highlighting their structured uncertainty—an attribute highly valued in games traditionally reliant on human partners. Xu et al.’s [173] exploration into LLMs within the social deduction game Werewolf demonstrates these AIs’ capability for strategic play, underscoring the necessity for advanced language understanding and social intuition. However, challenges such as limited input lengths and the impracticality of fine-tuning LLMs for specific games highlight current limitations.
Further critical examination by Sobieszek and Price [142] raises concerns over LLMs’ reliability in information delivery, emphasizing the risk of information pollution due to GPT-3’s occasionally untruthful responses. Research evaluating GPT models in social and moral contexts reveals mixed results, with models showing promise in single-player strategies but struggling in cooperative, ethical scenarios [2]. This suggests a need for further development in LLMs’ ability to navigate complex social interactions and ethical decision-making. Despite these challenges, the expressive capabilities of LLMs compared to earlier AI models mark a significant advancement in human-computer interaction within games.
5.4.2 LLMs for Content Generation in Games
Large Language Models have significantly impacted game development, from generating complex game levels as demonstrated by Todd et al.[150], to enriching RPGs with dynamic quest narratives through advancements shown by Vartinen et al. [156]. These models enhance game narratives and environment design, also facilitating interactive tutorials and support systems, a concept explored by [46] with Language Model Games. Despite their potential, challenges persist, including ensuring content quality and coherence, addressing ethical concerns, and maintaining player immersion. The instructRL framework by Hu et al. [65] offers insights into fine-tuning LLM integration for effective human-AI interaction in games. Advancing LLM applications in gaming requires addressing these hurdles through hybrid approaches and ethical guidelines.
5.5 Education
@@
Intelligent Tutoring Systems (ITSs), such as Carnegie Learning’s Mika platform, demonstrate AI’s ability to tailor education to individual needs, showing promising results like higher exam scores and reduced dropout rates [118, 155]. However, the opacity of AI’s decision-making has sparked debates over potential biases [168]. Despite these challenges, companies like Century Tech and Fishtree have successfully incorporated AI to enhance personalized learning experiences [76]. Virtual learning agents, including Georgia Tech’s chatbot Jill Watson, have been developed to supplement teaching by handling administrative queries or facilitating peer collaboration, albeit with risks of fostering parasocial relationships with digital avatars [40, 59]. Moreover, automated essay scoring systems offered by EdX and Pearson exemplify AI’s efficiency in grading, aligning closely with human grader assessments while saving time and costs [38]. Nonetheless, there remains skepticism about AI’s ability to fully grasp the subtleties of written expressions, highlighting the ongoing need for advancements in natural language processing.
@@
The use of LLMs in education is poised to personalize learning experiences further. Extance [42] explores the practical deployment of AI chatbots as interactive learning tools, demonstrating their potential to enhance the teaching and learning dynamic significantly. However, Milano et al. [141] raise critical points regarding the operational, ethical, and environmental considerations of LLMs in educational settings, emphasizing the need for sustainable and transparent practices. These insights offer a deeper look at AI’s role in education, balancing technological innovation with mindful consideration of its broader impacts [161, 61, 22].
5.6 Accessibility
@@
Assistive technologies powered by artificial intelligence (AI) are revolutionizing support for individuals with sensory and physical disabilities, offering innovative solutions for improved mobility and communication. A notable advancement is a deep learning-based navigation assistant for the visually impaired, which employs computer vision to facilitate obstacle detection and navigation with remarkable accuracy, as demonstrated in the work by [85]. Similarly, the development of communication aids for the deaf, utilizing audio-visual recognition algorithms, provides an automated solution that enhances daily interactions, a concept explored by [122]. These technologies exemplify the transformative potential of AI in creating more inclusive environments, with significant improvements in user independence and quality of life. Further contributions include a compact, low-cost visual aid system for blind individuals that integrates advanced image processing for obstacle avoidance and object detection [77], and an AI-enabled sign language recognition system that supports sentence-level communication for the hearing/speech impaired [166]. Additionally, the deployment of service robots to assist the elderly in shopping tasks demonstrates the practical applications of AI in facilitating daily activities [51].
@@
In addition to these technological innovations, recent studies highlight the opportunities and challenges presented by large language models in promoting accessibility. Taheri et al.[147] explore the potential of text-to-image generation models to revolutionize accessible art creation, demonstrating how generative text models can reduce typing effort and make visual art creation more accessible to people with diverse abilities. In contrast, Gadiraju et al.[49] shed light on the subtle yet harmful biases LLMs can perpetuate towards disabled communities, emphasizing the importance of incorporating disability-positive resources in training data to foster a more inclusive digital environment. These insights underscore the evolving landscape of AI-driven accessibility tools, stressing the need for continuous refinement and inclusivity in their development.
6 Conclusion

### v2

@@ 5. Applications
This section highlights the role of Human-AI (HAI) Teaming and Large Pre-trained Models (LPtM) in revolutionizing fields such as healthcare, autonomous vehicle domain, surveillance, gaming, education, and accessibility. It highlights the symbiosis between human expertise and AI’s analytical power, emphasizing the importance of ethical AI development and user-centric design in maximizing benefits and inclusivity within society.
5.1. Healthcare
@@
In the healthcare sector, recent researchers highlight the studies performed in enhancing patient diagnosis accuracy and fostering trust, emphasizing AI’s transformative influence in human-AI collaboration (Henry et al., 2022; Bienefeld et al., 2023; Memmert and Bittner, 2022). Advances such as the application of AI in breast cancer detection have markedly boosted diagnostic efficacy, sometimes even outperforming human experts, showcasing the remarkable potential of AI in healthcare (McKinney et al., 2020). These developments necessitate reevaluating AI’s role in patient care, addressing ethical considerations and practical challenges (Budd et al., 2021). Research focusing on healthcare practitioners’ views within the US highlights the critical factors of training, trustworthiness, and perceived AI risks, which are pivotal in its acceptance and application in clinical environments (Choudhury and Asan, 2022). Additionally, improvements in AI-powered medical image retrieval systems have been shown to augment both diagnostic capabilities and user trust, advocating for integrated systems that blend human insight with algorithmic accuracy (Cai et al., 2019). This partnership between AI technologies and healthcare professionals is vital for enhancing patient care outcomes and alleviating professional strain, albeit confronted by hurdles related to training and the accountability of AI systems.
@@
Integrating Large Pre-trained Models such as ChatGPT into healthcare has shown promise in enhancing medical education, patient communication, and clinical decision-making. Lyu et al. (Lyu et al., 2023) demonstrated ChatGPT’s ability to simplify radiology reports for patients and healthcare providers, highlighting the potential for improved clinical education despite noting areas for accuracy and specificity enhancements. Kung et al. (Kung et al., 2023) found that ChatGPT scored at or near the passing threshold on the USMLE, suggesting its capability to support medical education and hinting at future roles in clinical settings. Johnson et al. (Johnson et al., 2023) further validated ChatGPT’s usefulness by showing its generally accurate responses to a range of medical questions, though recognizing the need for further refinement.
Figure 7. An example of how multimodal foundation models and large language models can be used for enhancing generalization and robustness in autonomous driving, as studied by Wang et al. (Wang et al., 2023b). They develop pixel/patch-aligned feature descriptors and latent space simulation, enriched with language modality, suggesting potential for optimizing the training and debugging processes for end-to-end learning-based control. The RGB images in this figure are taken from the KITTI dataset (Geiger et al., 2012) for representation purposes.
5.2. Autonomous Vehicles
@@
Advancements in autonomous driving promise safer, AI-enhanced transportation, with significant milestones achieved in vehicle sensing and decision-making (Atakishiyev et al., 2021). Key to these advancements is understanding Human-AI collaboration, particularly during critical transitions between automated and manual driving modes, as explored by Lv et al. (Lv et al., 2021). This research highlights the importance of seamless transitions for safety, showing improvements in control and stability with a new hybrid driving model. Additionally, establishing trust in AI through transparent decision-making processes under uncertainty is crucial, as Jianlong and Chen (Zhou and Chen, 2019) discuss with a novel framework aimed at enhancing Human-AI collaboration. Fuchs et al. (Fuchs et al., 2023) further examine the managerial aspects of task delegation between humans and AI, proposing reinforcement learning to optimize collaboration and improve autonomous system performance, including driving.
@@
Building on foundational studies, recent explorations into large language models (LLMs) signal a transformative era in autonomous driving technologies. Yang et al.(Yang et al., 2024) underscore the importance of LLMs for interpreting user commands in complex scenarios, emphasizing the necessity of model quality and prompt design. Park et al. (Park et al., 2024) introduce VLAAD, a multi-modal LLM that demonstrates proficient interpretive capabilities across varied driving situations, showcasing the potential of natural language models in enhancing decision-making processes. Cui et al. (Cui et al., 2024) propose a framework that integrates LLMs for improved decision-making in autonomous vehicles, aiming for personalized assistance and efficient operation. Wang et al. (Wang et al., 2023a) discuss using LLMs as behavior planners to ensure safety in decision-making by interpreting driving scenes and adjusting trajectory plans based on safety constraints. More generally, Wen et al. (Wen et al., 2023) highlight the role of Visual Language Models (VLMs) like GPT-4V in scene understanding and decision-making, noting superior performance but also pointing out existing challenges. Additionally, Wang et al. (Wang et al., 2023b) describe an end-to-end autonomous driving system using multimodal foundation models to enhance adaptability and robustness in diverse environments, leveraging pixel-aligned features and language-based latent space simulation for improved data augmentation and policy debugging (Fig. 7). Together, these studies highlight the transformative potential of large foundation models in advancing autonomous vehicle technologies, from improving interpretive capabilities to enhancing decision-making processes and data generation.
5.3. Surveillance and Security
@@
The advancements in vision-based AI and IoT-connected cameras have heightened the potential for AI in cybersecurity (Pazho et al., 2023), including in Cyber Security Incident Response Teams (CSIRTs) (Killcrece et al., 2003). Determining the optimal level of AI autonomy in these settings is critical. A study by Hauptman et al. (Hauptman et al., 2023) explored this through a mixed-method approach, assessing AI’s role across the CSIRT cycle—preparation, identification, containment, eradication, and recovery. Surveying 103 cybersecurity professionals and conducting interviews with 22, the study found that higher AI autonomy is preferred during the identification phase, while lower autonomy suits high-risk phases better. This suggests that adaptive AI, with proper team training, can integrate effectively into CSIRTs, enhancing teamwork and response efficiency. Furthermore, the real-time update preference and successful integration of human and machine intelligence in crowd-powered camera sensing, as noted by Guo et al. (Guo et al., 2018), highlights the potential for enhancing cybersecurity measures through improved situational awareness and resource management.
@@
Moreover, the utilization of LLMs for video surveillance analysis, as proposed by Chen et al. (Chen et al., 2023), with the VideoLLM framework, demonstrates the potential for transferring LLMs’ understanding and reasoning capabilities to video understanding tasks, offering a unified approach for various video-based security challenges. Additionally, Jain’s (Jain, 2023) insights into integrating ChatGPT with video security underscore the gradually increasing impact of foundation models on physical security systems, enhancing operational efficiency and surfacing new insights for businesses.
5.4. Games
@@
5.4.1. LLM-Based AI as Players
Artificial intelligence has long been utilized in video games, from serving as basic adversaries in titles like Mario Kart (Switch, 2017) to mastering complex games such as chess, demonstrating AI’s algorithmic prowess. Yet, these AIs often fell short as complete substitutes for human players due to their lack of flexibility and limited language processing capabilities. The advent of LLMs has shifted this landscape, offering AIs that can engage with humans in more expressive ways, suitable across a broader spectrum of games.
Research into LLMs as game players, while in its nascent stages, has yielded intriguing insights. For instance, Frans (Frans, 2021) showcases "AI Charades" using GPT-2 to illustrate LLMs’ potential in games requiring deep language interaction, highlighting their structured uncertainty—an attribute highly valued in games traditionally reliant on human partners. Xu et al.’s (Xu et al., 2023) exploration into LLMs within the social deduction game Werewolf demonstrates these AIs’ capability for strategic play, underscoring the necessity for advanced language understanding and social intuition. However, challenges such as limited input lengths and the impracticality of fine-tuning LLMs for specific games highlight current limitations.
Further critical examination by Sobieszek and Price (Sobieszek and Price, 2022) raise concerns over LLMs’ reliability in information delivery, emphasizing the risk of information pollution due to GPT-3’s occasionally untruthful responses. Research evaluating GPT models in social and moral contexts reveals mixed results, with models showing promise in single-player strategies but struggling in cooperative, ethical scenarios (Akata et al., 2023). This suggests a need for further development in LLMs’ ability to navigate complex social interactions and ethical decision-making. Despite these challenges, the expressive capabilities of LLMs compared to earlier AI models mark a significant advancement in human-computer interaction within games.
5.4.2. LLMs for Content Generation in Games
Large Language Models have significantly impacted game development, from generating complex game levels as demonstrated by Todd et al.(Todd et al., 2023), to enriching RPGs with dynamic quest narratives through advancements shown by Vartinen et al. (Värtinen et al., 2022). These models enhance game narratives and environment design, also facilitating interactive tutorials and support systems, a concept explored by Frans (Frans, 2021) with Language Model Games. Despite their potential, challenges persist, including ensuring content quality and coherence, addressing ethical concerns, and maintaining player immersion. The instructRL framework by Hu and Sadigh (Hu and Sadigh, 2023) offer insights into fine-tuning LLM integration for effective human-AI interaction in games. Advancing LLM applications in gaming requires addressing these hurdles through hybrid approaches and ethical guidelines.
5.5. Education
@@
Intelligent Tutoring Systems (ITSs), such as Carnegie Learning’s Mika platform, demonstrate AI’s ability to tailor education to individual needs, showing promising results like higher exam scores and reduced dropout rates (Nwana, 1990; VanLehn, 2011). However, the opacity of AI’s decision-making has sparked debates over potential biases (Whitaker et al., 2013). Despite these challenges, companies like Century Tech and Fishtree have successfully incorporated AI to enhance personalized learning experiences (Kasinathan et al., 2017). Virtual learning agents, including Georgia Tech’s chatbot Jill Watson, have been developed to supplement teaching by handling administrative queries or facilitating peer collaboration, albeit with risks of fostering parasocial relationships with digital avatars (Eicher et al., 2018; Hartle and Kaczorowski, 2019). Moreover, automated essay scoring systems offered by EdX and Pearson exemplify AI’s efficiency in grading, aligning closely with human grader assessments while saving time and costs (Dikli, 2006). Nonetheless, there remains skepticism about AI’s ability to fully grasp the subtleties of written expressions, highlighting the ongoing need for advancements in natural language processing.
@@
The use of LLMs in education is poised to personalize learning experiences further. Extance (Extance, 2023) explores the practical deployment of AI chatbots as interactive learning tools, demonstrating their potential to enhance the teaching and learning dynamic significantly. However, Milano et al. (Silvia Milano, 2023) raise critical points regarding the operational, ethical, and environmental considerations of LLMs in educational settings, emphasizing the need for sustainable and transparent practices. These insights offer a deeper look at AI’s role in education, balancing technological innovation with mindful consideration of its broader impacts (Wang and Demszky, 2023; Hellas et al., 2023; Chang et al., 2023).
5.6. Accessibility
@@
Assistive technologies powered by artificial intelligence (AI) are revolutionizing support for individuals with sensory and physical disabilities, offering innovative solutions for improved mobility and communication. A notable advancement is a deep learning-based navigation assistant for the visually impaired, which employs computer vision to facilitate obstacle detection and navigation with remarkable accuracy, as demonstrated in the work by (Kumar and Jain, 2022). Similarly, the development of communication aids for the deaf, utilizing audio-visual recognition algorithms, provides an automated solution that enhances daily interactions, a concept explored by (Ozarkar et al., 2020). These technologies exemplify the transformative potential of AI in creating more inclusive environments, with significant improvements in user independence and quality of life. Further contributions include a compact, low-cost visual aid system for blind individuals that integrates advanced image processing for obstacle avoidance and object detection (Khan et al., 2020), and an AI-enabled sign language recognition system that supports sentence-level communication for the hearing/speech impaired (Wen et al., 2021). Additionally, the deployment of service robots to assist the elderly in shopping tasks demonstrates the practical applications of AI in facilitating daily activities (Ghazal et al., 2021).
@@
In addition to these technological innovations, recent studies highlight the opportunities and challenges presented by large language models in promoting accessibility. Taheri et al.(Taheri et al., 2023) explore the potential of text-to-image generation models to revolutionize accessible art creation, demonstrating how generative text models can reduce typing effort and make visual art creation more accessible to people with diverse abilities. In contrast, Gadiraju et al.(Gadiraju et al., 2023) shed light on the subtle yet harmful biases LLMs can perpetuate towards disabled communities, emphasizing the importance of incorporating disability-positive resources in training data to foster a more inclusive digital environment. These insights underline the evolving landscape of AI-driven accessibility tools, while also stressing the need for continuous refinement and inclusivity in their development.
6. Conclusion